EDITORIAL

Terry Goodman

We continue to receive feedback from readers concerning MJMS. This feedback is generally very positive with a variety of suggestions concerning issues such as the format of the journal, emphases of the articles, the reviewing process, number of issues per year, etc. Let me encourage you to communicate to us your questions, concerns, and suggestions for the journal. We want to be responsive to our subscribers and we welcome your input. Also, please share your copies of the journal with a colleague or two. As the number of subscribers to MJMS grows we will be able to provide more services to our readers.

While we have generally been pleased with the variety of manuscripts we have received, there are some topics/areas for which we would like to see more submissions. These include manuscripts focusing on topics in number theory, computer science, and mathematics education. With respect to the latter, manuscripts that focus on mathematics curriculum reform at the undergraduate or graduate levels would be welcomed.

Finally, we are considering producing a summer issue/monograph that would contain a compilation of all the problems that have appeared in the Problems section of our journal (over 100 problems by now). Whether we pursue this further will depend on the level of interest expressed by our readers. Please contact me with your thoughts concerning this possibility (see my mailing address and email address on inside back cover of this issue.) æ